# Spiking Neural Network Architecture Search: A Survey

### v1

@@ Abstract
This survey paper presents a comprehensive examination of Spiking Neural Network (SNN) architecture search (SNNaS) from a unique hardware/software co-design perspective. SNNs, inspired by biological neurons, have emerged as a promising approach to neuromorphic computing. They offer significant advantages in terms of power efficiency and real-time resource-constrained processing, making them ideal for edge computing and IoT applications. However, designing optimal SNN architectures poses significant challenges, due to their inherent complexity (e.g., with respect to training) and the interplay between hardware constraints and SNN models. We begin by providing an overview of SNNs, emphasizing their operational principles and key distinctions from traditional artificial neural networks (ANNs). We then provide a brief overview of the state of the art in NAS for ANNs, highlighting the challenges of directly applying these approaches to SNNs. We then survey the state-of-the-art in SNN-specific NAS approaches. Finally, we conclude with insights into future research directions for SNN research, emphasizing the potential of hardware/software co-design in unlocking the full capabilities of SNNs. This survey aims to serve as a valuable resource for researchers and practitioners in the field, offering a holistic view of SNNaS and underscoring the importance of a co-design approach to harness the true potential of neuromorphic computing.
@@ I Introduction
SNNs offer several potential advantages over ANNs. Spike timing in SNNs captures temporal patterns better than static ANNs, mirroring real-world signals that are inherently temporal, such as audio, visual, and sensory data [1, 2]. Their sparsity ensures energy efficiency, making them ideal for constrained environments [3]. SNNs’ ability to mimick biological neurons enables the potential to derive insight into brain function, advancing neuroscience and AI while boosting network fault tolerance [4, 5]. The temporal nature and unique structure of SNNs make them efficient, versatile, and closer to natural neural processes than traditional artificial networks [6].
@@
Given these challenges, Spiking Neural Network Architecture Search (SNNaS) has emerged as a critical research direction [13, 14, 7, 15, 16, 8, 17, 18, 19, 20, 12, 21, 22, 23, 24, 25, 26, 27, 28]. SNNaS aims to automate the discovery of optimal network architectures by simultaneously optimizing multiple objectives, including inference accuracy, processing speed, and energy consumption. Importantly, this optimization process must consider both software and hardware constraints, particularly when targeting deployment on neuromorphic hardware [29].
@@
This survey paper provides a comprehensive analysis of SNNaS research through the lens of hardware-software co-design. We begin by establishing fundamental SNN concepts, including spike encoding/decoding mechanisms, neuron models, and training methodologies, while highlighting key distinctions from traditional ANNs. Building on this foundation, we examine how Neural Architecture Search (NAS) approaches from the ANN domain must be adapted for the unique characteristics of SNNs. The survey’s core contribution lies in its systematic exploration of SNNaS methodologies, encompassing search space design, optimization algorithms, acceleration strategies, and performance estimation techniques, with particular emphasis on hardware constraint integration. Finally, looking forward, we identify promising research directions, including surrogate gradient optimization, Bayesian approaches, hybrid methodologies, and emerging neuromorphic hardware platforms. Through this analysis, we aim to illuminate both the challenges and opportunities in developing efficient SNN architectures for next-generation AI systems.
A chevron flowchart showing the Neural Architecture Search (NAS) pipeline for spiking neural networks, read left-to-right as a series of chevron-shaped blocks. The first block is the Search Space and has three child bubbles: Layer Configurations, Neuron Models, and Hyperparameters. The second block is the Search Strategy, which has three child bubbles that list example search methods: Evolutionary Algorithms, Gradient-Based Search, and Reinforcement Learning. The next block, ‘Performance Evaluation,’ has three child bubbles: Accuracy, Latency, and Energy. The final block is ‘Architecture Selection.’ The diagram illustrates the pipeline, which involves defining the search space, selecting and executing a search strategy, evaluating candidate architectures on accuracy, latency, and energy, and finally selecting the architecture.
@@
Figure 1: Overview of the Neural Architecture Search (NAS) workflow for Spiking Neural Networks: From defining the search space (e.g., layer configurations, neuron models, and hyperparameters) to employing search strategies (e.g., evolutionary algorithms, gradient-based search, and reinforcement learning, etc.) for performance evaluation (e.g., accuracy, latency, and energy) and, finally, optimal architecture selection.
The remainder of this paper is organized as follows: Section II introduces the fundamental concepts of SNNs, including their operational principles, neuron models, spike encoding mechanisms, and training methodologies. Section III examines Neural Architecture Search, beginning with the unique challenges of applying NAS to SNNs, followed by an extensive analysis of search spaces, search strategies, and evaluation methods. We also discuss various acceleration techniques and hardware considerations in this section. Section IV explores current challenges and future research directions, including scalability issues, energy efficiency optimization, and the development of more sophisticated training algorithms. Finally, Section V concludes by summarizing the key insights and emphasizing the importance of hardware/software co-design in advancing SNNaS research.
@@ II-A1 Neuron Models
SNNs employ diverse neuron models to emulate biological neurons, presenting different trade-offs for architecture search. For example, the Leaky Integrate-and-Fire (LIF) model [32, 30] is a widely used model that describes neuron behavior through membrane potential dynamics and spike generation, balancing computational efficiency with biological plausibility. The Hodgkin-Huxley model [33], on the other hand, offers high biological accuracy but demands significant computational resources due to its granular modeling of ion channel kinetics and membrane potentials. The Izhikevich model [34] strikes a balance between the biological plausibility of the Hodgkin-Huxley model and the computational efficiency of the LIF model [35]. It abstracts the low-level details present in the Hodgkin-Huxley models while maintaining key dynamic properties.
Each model presents unique advantages, allowing researchers to select the most appropriate approach for their specific SNN applications, often balancing the trade-off between biological accuracy and computational efficiency. For example, while the precision of the Hodgkin-Huxley model makes it ideal for studies requiring biophysical detail, its computational intensity can limit scalability. In contrast, the Izhikevich model can replicate a wide range of spike behaviors with significantly lower computational demands. However, this efficiency comes at the expense of detailed biophysical insights, which may be critical for applications requiring fine-grained neuronal dynamics.
@@ II-A2 Spike Encoding
Various encoding algorithms transform input data into spike trains, each suited to specific applications and hardware constraints [36, 37]. Rate-based encoding, for example, uses the firing frequency to represent signal magnitude [38]. Temporal encoding methods focus on the timing of spikes, offering different strategies: time-to-first-spike emits earlier spikes for larger signals [39]; phase coding periodically modulates spike weights [40]; and burst encoding utilizes inter-spike intervals and bursts to convey information [41].
Spatio-temporal encoding captures both spike timing and spatial distribution across neurons or layers, enhancing the representation of neural dynamics and supporting complex pattern learning [42]. Population encoding further extends this by involving neuron groups to represent single inputs, with each neuron responding to distinct signal features to enable robust and distributed representations.
[…]
@@ II-A3 Training Approaches
Synaptic plasticity underpins learning in biological brains, with connection strengths changing over time [44]. SNNs aim to mimic this but face unique training challenges. The non-differentiable spike function complicates traditional backpropagation methods [45]. Furthermore, training deep SNNs with multi-layer learning is particularly challenging, often resulting in learning being confined to a single layer in many SNNs [46, 47, 48]. While progress has been made with two-layer [49, 50, 51] and recurrent SNNs [52, 53], training deeper networks remains difficult. Hidden layers are crucial for tackling complex real-world problems, making this a key hurdle to overcome [54].
@@
Three primary methods are used for training SNNs and influence architecture design: surrogate gradients, ANN-to-SNN conversion, and spike-timing-dependent plasticity (STDP).
@@
Surrogate gradients enable SNN training by approximating non-differentiable spike functions, allowing backpropagation while preserving temporal dynamics [54]. Building on initial work [55, 56], various backpropagation through time (BPTT) methods have demonstrated effectiveness in image classification [54, 57, 58]. While these approaches achieve high accuracy with low latency, their resource intensity limits application to simpler architectures [59]. The optimal choice of surrogate function must balance accuracy and implementation efficiency. Although software implementations offer flexibility [54], hardware considerations impact spiking activity sparsity and implementation efficiency [60].
@@
ANN-to-SNN Conversion (or Shadow Training)transforms ANNs into SNNs by training a ‘shadow’ ANN, then replacing analog neurons with spiking neurons, while retaining weights and reconfiguring neurons to mimic original network. This bypasses the issue of nondifferentiable spiking activity and reduces training overhead [61]. The spike rate or latency codes approximate the ANN’s behavior, but this conversion poses challenges such as requiring lengthy simulations that can negate SNNs’ power and speed advantages [62]. Moreover, the approximation process may lead to performance degradation, especially as latency is decreased [63, 64]. Additionally, since ANNs are often trained on static datasets like MNIST and CIFAR-10, the converted SNNs may underutilize temporal dynamics, limiting their potential advantages [3].
Spike-Timing Dependent Plasticity (STDP) [65] adjusts synapse strength based on spike timing, acting as an unsupervised learning rule for SNNs. Hebbian STDP strengthens synapses when pre-synaptic neurons fire before post-synaptic neurons and weakens them otherwise, while anti-Hebbian STDP inverts this pattern. The proximity of spike pairs determines weight adjustments, making STDP sensitive to order and timing. While STDP naturally adapts to temporal pattern learning, its unsupervised nature complicates network control and does not guarantee optimal performance, especially in complex tasks [66]. Its effectiveness diminishes in deeper networks, limiting its scalability in multilayered SNNs [67], and STDP-trained networks often underperform other learning algorithms [68].
@@ II-B Challenges of SNNs
Table I illustrates the fundamental paradox of SNNs: while offering compelling advantages like energy efficiency and temporal processing, each benefit introduces corresponding challenges that must be carefully balanced in the design process. First, as previously alluded to, their non-differentiable spiking mechanism complicates the training process, making standard techniques like backpropagation unsuitable without adaptations [3]. While techniques like surrogate gradients have become a workhorse for training SNNs, they remain less mature and efficient compared to those used in traditional ANNs.
SNNs can be computationally costly due to their temporal dynamics. Unlike ANNs, which compute activations in a single pass, SNNs require the simulation of multiple time steps to capture spike-based neuronal dynamics, increasing both computational cost and latency [69]. This complexity becomes more pronounced when scaling SNNs to large networks or datasets.
@@
Key issues such as identifying optimal SNN designs, understanding the impact of spike encoding and decoding, and scaling SNNs to handle larger and more complex datasets remain open challenges [71]. Addressing these challenges requires systematic exploration of the design space to develop efficient and scalable architectures. Given the computational constraints and temporal dynamics of SNNs, hardware-software co-design plays a critical role in ensuring that these architectures perform effectively on neuromorphic hardware [37]. This necessitates the studies on Neural Architecture Search (NAS), which has emerged as a promising approach for automating the discovery of optimal SNN designs. By integrating NAS with hardware-aware techniques, researchers can identify architectures that balance accuracy, efficiency, and resource constraints, unlocking the full potential of SNNs for real-world applications.
@@
Biological Plausibility Mimics biological neurons, offering insights into neuroscience and fault-tolerant systems. Training non-differentiable spike-based mechanisms remains challenging.
@@
Training Surrogate gradient and conversion techniques enable learning despite spiking non-differentiability. Deep multi-layer SNNs remain difficult to train effectively, needing further research.
Hardware Integration Efficiently integrates with neuromorphic chips (e.g., Loihi) for improved hardware utilization. Neuromorphic hardware often supports only limited neuron and synapse model diversity.
@@ III Neural Architecture Search
NAS automates neural network design by treating it as an optimization problem, seeking architectures that maximize performance—typically measured as validation accuracy—while potentially considering other objective functions like computational efficiency. Instead of manually designing networks, NAS employs optimization techniques to identify optimal structures for a given task. An early NAS approach by Stanley et al. [72] used evolutionary methods, inspiring numerous refinements. Although these advances have benefited ANNs, SNNs pose distinct challenges due to their event-driven processing and temporal dynamics, necessitating specialized architectural considerations. To fully exploit SNNs’ potential, NAS methods must be adapted accordingly. This section explores why traditional NAS techniques designed for ANNs may be ineffective for SNNs and examines existing NAS approaches for SNNs, focusing on the three core components: search space, search strategy, and evaluation [73] (Figure 2).
A diagram of an iterative Neural Architecture Search loop. From left to right, three rectangular boxes read ‘Search Space,’ ‘Search Strategy,’ and ‘Evaluation Method.’ A straight arrow connects Search Space to Search Strategy. From Search Strategy to Evaluation Method, a curved arrow across the top is labeled ‘Searched Architectures.’ From Evaluation Method back to Search Strategy, a curved arrow along the bottom is labeled ‘Performance Feedback.’ The arrows form a closed loop, indicating that candidate architectures are evaluated and their results are fed back to guide further search.
Figure 2: Neural Architecture Search has three important components: search space, search strategy, and evaluation [73].

### v2

@@ Abstract
This survey paper presents a comprehensive examination of Spiking Neural Network (SNN) architecture search (SNNaS) from a unique hardware/software co-design perspective. SNNs, inspired by biological neurons, have emerged as a promising approach to neuromorphic computing. They offer significant advantages in terms of power efficiency and real-time resource-constrained processing, making them ideal for edge computing and IoT applications. However, designing optimal SNN architectures poses significant challenges, due to their inherent complexity (e.g., with respect to training) and the interplay between hardware constraints and SNN models. We begin by providing an overview of SNNs, emphasizing their operational principles and key distinctions from traditional artificial neural networks (ANNs). We then provide a brief overview of the state of the art in NAS for ANNs, highlighting the challenges of directly applying these approaches to SNNs. We then survey the state of the art in SNN-specific NAS approaches. Finally, we conclude with insights into future research directions for SNN research, emphasizing the potential of hardware/software co-design in unlocking the full capabilities of SNNs. This survey aims to serve as a valuable resource for researchers and practitioners in the field, offering a holistic view of SNNaS and underscoring the importance of a co-design approach to harness the true potential of neuromorphic computing.
@@ I Introduction
Given these challenges, Spiking Neural Network Architecture Search (SNNaS) has emerged as a critical research direction [100, 97, 29, 99, 68, 30, 132, 167, 84, 114, 112, 111, 105, 170, 82, 102, 5, 160]. As illustrated in Figure 1, SNNaS methods comprise three core components: a search space defining architectural choices (e.g., layer configurations, neuron models, hyperparameters), a search strategy for exploring this space (e.g., evolutionary algorithms, gradient-based methods, reinforcement learning), and a performance evaluation mechanism that guides architecture selection based on metrics such as accuracy, latency, and energy consumption. SNNaS methods focus on automating the discovery of optimal SNN architectures by simultaneously optimizing multiple, often competing objectives, including inference accuracy, processing speed, and energy consumption. Importantly, this optimization process must consider both software and hardware constraints, particularly when targeting deployment on neuromorphic hardware [149].
Despite the growing body of literature on spiking neural networks, most existing surveys focus on foundational aspects such as neuron models, training methodologies, learning rules, hardware implementations, or application domains. These works provide essential background on how SNNs operate and how they can be trained or deployed, but the problem of automating architectural design choices—long recognized as critical for ANN performance and efficiency—has only recently begun to receive focused attention in the SNN domain.
@@
This survey paper provides a comprehensive analysis of SNNaS research through the lens of hardware-software co-design. We begin by establishing fundamental SNN concepts, including spike encoding/decoding mechanisms, neuron models, and training methodologies, while highlighting key distinctions from traditional ANNs. Building on this foundation, we examine how Neural Architecture Search (NAS) approaches from the ANN domain must be adapted for the unique characteristics of SNNs. The survey’s core contribution lies in its systematic exploration of SNNaS methodologies, encompassing search space design, search algorithms, acceleration strategies, and performance estimation techniques, with particular emphasis on integrating hardware constraints. Finally, looking forward, we identify promising research directions, including surrogate gradient optimization, Bayesian approaches, hybrid methodologies, and emerging neuromorphic hardware platforms. Through this analysis, we aim to illuminate both the challenges and opportunities in developing efficient SNN architectures for next-generation AI systems.
@@
Figure 1: Overview of the NAS workflow for SNNs. The pipeline defines the three NAS components within an SNN-specific roadmap: 1) search space, 2) search strategy, and 3) evaluation [46]. The search space defines architectural choices such as layer configurations, neuron models, and hyperparameters; search strategies explore this space (e.g., evolutionary algorithms, gradient-based methods, and reinforcement learning); and candidate architectures are evaluated using performance metrics including accuracy, latency, and energy to guide final architecture selection.
In summary, the key insights of this survey are as follows:
  1. 1.
    Hardware-software co-design is essential, not optional. SNN-specific temporal dynamics and event-driven computation create fundamental mismatches with ANN-centric NAS methods, requiring integrated consideration of algorithmic and hardware constraints from the outset.
  2. 2.
    Search space design critically determines SNNaS success. The field has evolved from naively replacing ReLu with LIF neurons toward SNN-specific representations (e.g., cell-based directed acyclic graphs DAGs, hierarchical structures, and biologically-inspired motifs) that better capture spiking computation.
  3. 3.
    Temporal dynamics remain underexploited. Despite being a defining advantage of SNNs, many of the surveyed SNNaS methods adapt ANN-centric approaches that overlook time-related factors such as spike timing, timestep optimization, and temporal attention.
  4. 4.
    Training-free evaluation strategies are transforming efficiency. Zero-shot metrics, surrogate models, and one-shot supernets have reduced search costs by orders of magnitude, making SNNaS practical for resource-constrained settings.
  5. 5.
    Multi-objective optimization is inherent to SNNaS. Unlike traditional NAS, SNNaS must simultaneously balance accuracy, energy, latency, memory, and spike sparsity. This requirement aligns naturally with evolutionary and Bayesian approaches.
  6. 6.
    Co-exploration outperforms sequential optimization. Frameworks that jointly search neural architectures and hardware configurations consistently achieve better energy-delay products than methods optimizing these dimensions separately.
@@ II-A1 Neuron Models
SNNs employ diverse neuron models that trade biological plausibility for computational efficiency. The Leaky Integrate-and-Fire (LIF) model [69, 49] strikes a practical balance between these factors and therefore dominates SNNaS research. In contrast, the Hodgkin–Huxley model [57] offers high biophysical fidelity but incurs substantial computational cost. The Izhikevich model [61] occupies an intermediate position, reproducing a wide range of spiking behaviors with relatively low computational overhead [145], albeit without detailed biophysical interpretability. Crucially, this trade-off between fidelity and simplicity propagates beyond modeling, directly affecting architecture search space complexity and hardware implementation efficiency.
@@ II-A2 Spike Encoding
Various encoding algorithms transform input data into spike trains, each suited to specific applications and hardware constraints [10, 9]. Rate-based encoding, for example, uses the firing frequency to represent signal magnitude [52]. Temporal encoding methods focus on the timing of spikes, offering different strategies: time-to-first-spike emits earlier spikes for larger signals [107]; phase coding periodically modulates spike weights [67]; burst encoding utilizes inter-spike intervals and bursts to convey information [60]; and spatio-temporal encoding captures both spike timing and spatial distribution across neurons or layers [131].
The choice of encoding scheme significantly influences the efficiency of information transfer and the performance of both hardware and software. Spatio-temporal encoding offers high efficiency for tasks requiring dynamic interaction between neurons [131], while temporal pattern coding can reduce power consumption in computationally demanding tasks [122]. These trade-offs highlight the importance of aligning encoding strategies with the specific requirements of both the application and the underlying hardware.
@@ II-A3 Training Approaches
Synaptic plasticity underpins learning in biological brains, with connection strengths changing over time [1]. SNNs aim to mimic this but face unique training challenges. The non-differentiable spike function complicates traditional backpropagation methods [136]. Furthermore, training deep SNNs with multi-layer learning is particularly challenging, often resulting in learning being confined to a single layer in many SNNs [92, 19, 148].
@@
Four primary methods used for training SNNs influence architecture design: surrogate gradients, ANN-to-SNN conversion, spike-timing-dependent plasticity (STDP), and evolutionary optimization (EO).
@@
Surrogate gradients enable SNN training by approximating non-differentiable spike functions, allowing backpropagation while preserving temporal dynamics [98]. While backpropagation through time (BPTT) methods have demonstrated high accuracy with low latency on image classification tasks [98, 22, 164], their resource intensity limits application to simpler architectures [118]. The optimal choice of surrogate function must balance accuracy and implementation efficiency. Although software implementations offer flexibility [98], hardware considerations impact spiking activity sparsity and implementation efficiency [7].
@@
ANN-to-SNN Conversion transforms ANNs into SNNs by training a ‘shadow’ ANN, then replacing analog neurons with spiking neurons, retaining ANN weights, and reconfiguring neurons to mimic the original network. This bypasses the issue of non-differentiable spiking activity and reduces training overhead [158]. The spike rate or latency codes approximate the ANN’s behavior, but this conversion can require lengthy simulations that can negate SNNs’ power and speed advantages [123]. Moreover, the approximation process may degrade performance, especially as latency decreases [39, 56]. Additionally, since ANNs are often trained on static datasets like MNIST and CIFAR-10, the converted SNNs may underutilize temporal dynamics, limiting their potential advantages [38]. As a result, ANN-to-SNN conversion methods primarily constrain architecture design rather than discover SNN-native structures, and are therefore best interpreted as baselines or transitional approaches rather than fully spiking-native SNNaS methods
Spike-Timing Dependent Plasticity (STDP) [20] adjusts synapse strength based on spike timing as an unsupervised learning rule. While naturally adapted to temporal patterns, its effectiveness diminishes in deeper networks [154], and STDP-trained networks often underperform other algorithms [79].
Evolutionary Optimization for Neuromorphic Systems (EONS) [127] uses evolutionary optimization (EO), a gradient-free method, to jointly evolve network topology and neuron/synapse parameters (e.g., weights, thresholds, synaptic delays). Because EO does not rely on gradient information, it bypasses the challenges posed by non-differentiable spike activation functions. As a highly versatile approach, it enables SNN design for classification, control, and reservoir generation under neuromorphic hardware constraints [128]. The disadvantage is that EO methods are often computationally expensive, requiring substantial compute to reach convergence.
@@ II-B Challenges of SNNs
Table I illustrates the fundamental paradox of SNNs: while offering compelling advantages like energy efficiency and temporal processing, each benefit introduces corresponding challenges that must be carefully balanced in the design process. First, as previously alluded to, their non-differentiable spiking mechanism complicates the training process, making standard techniques like backpropagation unsuitable without adaptations [38]. While techniques like surrogate gradients have become a workhorse for training SNNs, they remain less mature and efficient compared to those used in traditional ANNs. Furthermore, many traditional training approaches fail to account for the unique restrictions of neuromorphic hardware, such as physical connectivity limitations, limited synaptic weight resolution, and the presence of operational noise. These discrepancies can lead to significant performance degradation when a trained network is eventually mapped onto a physical system.
SNNs can be computationally costly due to their temporal dynamics. Unlike ANNs, which compute activations in a single pass, SNNs require multiple time steps to capture spike-based neuronal dynamics, increasing both computational cost and latency [152]. This complexity is further exacerbated during the training phase. For instance, evolutionary approaches like EONS [128] can require a large number of epochs and a large population size to converge, necessitating substantial compute resources and parallelization. Additionally, a "hardware/software boundary" bottleneck may arise if a software simulator is unavailable, due to the high communication overhead required to configure and evaluate diverse network structures on neuromorphic hardware [127]. However, these challenges can be partially mitigated through multi-objective optimization, which allows for the evolution of non-traditional, highly simplified network structures that use much fewer neurons and synapses than standard architectures, thereby improving area and power efficiency [127].
@@
Key issues such as identifying optimal SNN designs, understanding the impact of spike encoding and decoding, and scaling SNNs to handle larger and more complex datasets remain open challenges [23]. Addressing these challenges requires systematic exploration of the design space to develop efficient and scalable architectures. Given the computational constraints and temporal dynamics of SNNs, hardware-software co-design plays a critical role in ensuring that these architectures perform effectively on neuromorphic hardware [9]. This necessitates studies on NAS, which has emerged as a promising approach for automating the discovery of optimal SNN designs. By integrating NAS with hardware-aware techniques, researchers can identify architectures that balance accuracy, efficiency, and resource constraints, unlocking the full potential of SNNs for real-world applications.
@@
Biological Plausibility Mimics biological neurons, offering insights into neuroscience and fault-tolerant systems. Training non-differentiable spike-based mechanisms remains challenging for gradient-based methods [127].
@@
Training Surrogate gradient, conversion, and EO techniques enable learning despite spiking non-differentiability. Deep multi-layer SNNs remain difficult to train effectively, needing further research.
Hardware Integration Multi-objective optimization produces smaller, power-efficient networks tailored to specific hardware constraints for better integration with neuromorphic chips (e.g., Loihi [84]) and improved hardware utilization. Neuromorphic hardware often supports only limited neuron and synapse model diversity. High communication overhead at the "hardware/software boundary" can bottleneck training if a simulator is unavailable [127].
@@ III Neural Architecture Search
NAS automates the tedious manual process of neural network design by employing optimization techniques that identify and rank optimal architectures for a given task based on a fitness metric (e.g., accuracy or computational efficiency). This search for the optimal configuration can be accelerated by approximating the ranking without fully training every candidate network. A misranking can occur when approximate evaluation produces a different ordering than full training with temporal execution.
An early NAS approach by Stanley et al. [140] used evolutionary methods, inspiring numerous refinements. Although these advances have benefited ANNs, SNNs pose distinct challenges due to their event-driven processing and temporal dynamics, necessitating specialized architectural considerations. To fully exploit SNNs’ potential, NAS methods must be adapted accordingly. This section explores why traditional NAS techniques designed for ANNs may be ineffective for SNNs and examines existing NAS approaches for SNNs, focusing on the three core components: search space, search strategy, and evaluation [46] (Figure 1).